# Data-Driven Distributed Optimization via Aggregative Tracking and Deep-Learning

### v2

@@ VI Conclusions
We proposed DELTA, a data-driven distributed algorithm for aggregative optimization in unknown scenarios. DELTA combines an optimization-oriented part with, a tracking mechanism, and a learning strategy. The latter takes on local neural networks elaborating feedback of the local costs evaluated around the current solution estimates to approximate their gradients. We analyzed DELTA by using tools from system theory based on timescale separation and averaging theory. More in detail, we theoretically proved that the trajectories of the obtained interconnected system linearly converge in a neighborhood of the optimal solution. The size of such a neighborhood depends on the given accuracy capabilities of the neural networks. Some numerical simulations confirmed our theoretical findings.
@@ A-A Preliminaries on Averaging Theory
We report [38, Theorem 2.2.2], which is a useful result in the context of averaging theory. Consider the time-varying system
@@
$f_{\text{\tiny AV}}(\chi)\coloneq\lim_{T\to\infty}\frac{1}{T}\sum_{\tau=\bar{t}+1}^{\bar{t}+T}f(\chi,0,\tau),$ (40)
@@
[38, Theorem 2.2.3] Consider system (39) and let Assumptions 789, and 10 hold. If there exists $\bar{\gamma}_{1}>0$ such that, for all $\gamma\in(0,\bar{\gamma}_{1})$, the origin is an exponentially stable equilibrium point of the averaged system (41), then there exists $\bar{\gamma}\in(0,\bar{\gamma}_{1})$ such that, for all $\gamma\in(0,\bar{\gamma})$, the origin is an exponentially stable equilibrium point of system (39). $\blacksquare$
@@ B-B Proof of Lemma 1
We point out that Assumption (3) implies that $R^{\top}AR$ is Schur. Hence, for all $q,>0$, there exist $\Pi_{1}=\Pi_{1}^{\top}>0$ and $\Pi_{2}=\Pi_{2}^{\top}>0$ solving the discrete-time Lyapunov equations
@@
$\left(R^{\top}AR\right)^{\top}\Pi_{2}\left(R^{\top}AR\right)-\Pi_{2}$ $=I_{(N-1)d},$ (43b)
@@
Being $U$ quadratic, the conditions (24a) and (24c) are satisfied. Then, to conclude the proof, we only need to show that (24b) is verified too. To this end, we study the increment of $U$ along the trajectories of system (23) obtaining
@@
$\left\lVert\Delta\hat{G}_{2}(\tilde{\mathrm{x}}\!+\!x^{\star}\!,\tilde{\mathrm{w}}_{\perp},\tilde{\theta}\!+\!\theta^{\star}\!)\right\rVert\leq L_{\sigma}\left\lVert R\tilde{\mathrm{w}}_{\perp}\right\rVert$, (45)
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}-\!q\left\lVert\zeta_{1}\right\rVert^{2}\!-\!\left\lVert\zeta_{2}\right\rVert^{2}\!+\!2\left\lVert\zeta_{2}\right\rVert\left\lVert R^{\top}AR\Pi_{2}RA\right\rVert\!\left\lVert\Delta\hat{G}_{2}(\mathrm{x},\zeta_{1},\tilde{\theta})\right\rVert+\left\lVert(R^{\top}M)\!^{\top}\!\Pi_{2}(R^{\top}M)\right\rVert\left\lVert\Delta\hat{G}_{2}(\mathrm{x},\zeta_{1},\tilde{\theta})\right\rVert^{2}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\!-q\left\lVert\zeta_{1}\right\rVert^{2}-\left\lVert\zeta_{2}\right\rVert^{2}+2L_{\sigma}\left\lVert R^{\top}AR\Pi_{2}RA\right\rVert\left\lVert R\zeta_{1}\right\rVert\left\lVert\zeta_{2}\right\rVert+\left\lVert(R^{\top}M)\!^{\top}\!\Pi_{2}(R^{\top}M)\right\rVert L_{\sigma}^{2}\left\lVert R\zeta_{1}\right\rVert^{2},$ (46)
@@
$\begin{cases}q&>\kappa_{1}\\ q&>\kappa_{1}+\kappa_{12}^{2}.\end{cases}$ (48)
By setting set $q>\bar{q}_{1}\coloneq\kappa_{1}+\kappa_{12}^{2}$, we have $P_{U}>0$ and, thus, we bound (47) as
@@ C-C Proof of Lemma 2
$\Delta W_{\theta}(\tilde{\theta}_{\text{\tiny AV}},\tilde{\mathrm{x}}_{\text{\tiny AV}})$ $\coloneq\left\lVert\tilde{\theta}_{\text{\tiny AV}}-\gamma\nabla_{2}\tilde{\ell}_{\mathcal{K}}(\tilde{\mathrm{x}}_{\text{\tiny AV}},\tilde{\theta}_{\text{\tiny AV}})\right\rVert^{2}-\left\lVert\tilde{\theta}_{\text{\tiny AV}}\right\rVert^{2}$
$\Delta W_{\mathrm{x}}(\tilde{\theta}_{\text{\tiny AV}},\tilde{\mathrm{x}}_{\text{\tiny AV}})$ $\coloneq\left\lVert\tilde{\mathrm{x}}_{\text{\tiny AV}}-\gamma\hat{G}^{d}(\tilde{\mathrm{x}}_{\text{\tiny AV}},\tilde{\theta}_{\text{\tiny AV}})\right\rVert^{2}-\left\lVert\tilde{\mathrm{x}}_{\text{\tiny AV}}\right\rVert^{2}.$
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}\!-\gamma\frac{2\mu_{\ell}L_{\ell}}{\mu_{\ell}\!+\!L_{\ell}}\!\left\lVert\tilde{\theta}_{\text{\tiny AV}}\right\rVert^{2}\!\!-\!\gamma\!\left(\frac{2}{\mu_{\ell}\!+\!L_{\ell}}\!-\!\gamma\right)\!\left\lVert\nabla_{2}\tilde{\ell}_{\mathcal{K}}(\tilde{\mathrm{x}}_{\text{\tiny AV}},\tilde{\theta}_{\text{\tiny AV}})\right\rVert^{2}\!\!$ (52)
where in $(a)$ we add $\nabla_{2}\tilde{\ell}_{\mathcal{K}}(\tilde{\mathrm{x}}_{\text{\tiny AV}},0)=0$, then we use the $\mu_{\ell}$-strong convexity of $\tilde{\ell}_{\mathcal{K}}$ in $\Theta$ and the $L_{\ell}$-Lipschitz continuity of $\nabla_{2}\tilde{\ell}_{\mathcal{K}}$ (cf. Assumption 6). Similarly, we expand the square norm in $\Delta W_{\mathrm{x}}$ and obtain
@@
$\!\stackrel{{\scriptstyle(a)}}{{=}}-2\gamma\nabla f_{\sigma}(\tilde{\mathrm{x}}_{\text{\tiny AV}}+x^{\star})^{\top}\tilde{\mathrm{x}}_{\text{\tiny AV}}-2\gamma(\hat{G}^{d}(\tilde{\mathrm{x}}_{\text{\tiny AV}},\tilde{\theta}_{\text{\tiny AV}})\!-\!\hat{G}^{d}(\tilde{\mathrm{x}}_{\text{\tiny AV}},0))^{\top}\!\tilde{\mathrm{x}}_{\text{\tiny AV}}$
@@
$\!\stackrel{{\scriptstyle(b)}}{{=}}-2\gamma\nabla f_{\sigma}(\tilde{\mathrm{x}}_{\text{\tiny AV}}+x^{\star})^{\top}\tilde{\mathrm{x}}_{\text{\tiny AV}}+2\gamma^{2}\left\lVert\hat{G}^{d}(\tilde{\mathrm{x}}_{\text{\tiny AV}},0)\right\rVert^{2}\!\!+2\gamma\!\left\lVert\hat{G}^{d}(\tilde{\mathrm{x}}_{\text{\tiny AV}},\tilde{\theta}_{\text{\tiny AV}})\!-\!\hat{G}^{d}(\tilde{\mathrm{x}}_{\text{\tiny AV}},0)\right\rVert\left\lVert\tilde{\mathrm{x}}_{\text{\tiny AV}}\right\rVert$
@@
where in $(a)$ we add $\pm\nabla f_{\sigma}(\tilde{\mathrm{x}}_{\text{\tiny AV}}+x^{\star})\!=\!\hat{G}^{d}(\tilde{\mathrm{x}}_{\text{\tiny AV}},0)$, and $(b)$ uses the triangle and Cauchy-Schwarz inequalities and $\nabla f_{\sigma}(x^{\star})\!=\!\hat{G}^{d}(0,0)\!=\!0$. By Lipschitz continuity of $\nabla_{1}\hat{f}_{i}$, $\nabla_{2}\hat{f}_{i}$, and $\nabla\phi_{i}$ (cf. Assumptions 2 and 4), there exists $L_{\theta}\!>\!0$ such that
$\left\lVert\hat{G}^{d}(\tilde{\mathrm{x}}_{\text{\tiny AV}},\tilde{\theta}^{k}_{\text{\tiny AV}})-\hat{G}^{d}(\tilde{\mathrm{x}}_{\text{\tiny AV}},\bar{\theta}_{\text{\tiny AV}})\right\rVert\leq L_{\theta}\left\lVert\tilde{\theta}_{\text{\tiny AV}}-\bar{\theta}_{\text{\tiny AV}}\right\rVert,$ (54)
for all $\tilde{\mathrm{x}}_{\text{\tiny AV}}\in\mathbb{R}^{n}$ and $\tilde{\theta}_{\text{\tiny AV}},\bar{\theta}_{\text{\tiny AV}}\in\mathbb{R}^{m}$. Using (54), the $\mu_{f}$-strong convexity, and $L_{x}$-Lipschitz continuity of $f_{\sigma}$, we bound (53) as
$\Delta W_{\mathrm{x}}(\tilde{\theta}_{\text{\tiny AV}},\tilde{\mathrm{x}}_{\text{\tiny AV}})$ $\leq-\gamma\frac{2\mu_{f}L_{x}}{\mu_{f}+L_{x}}\left\lVert\tilde{\mathrm{x}}_{\text{\tiny AV}}\right\rVert^{2}-\gamma\left(\frac{2}{\mu_{f}+L_{x}}-\gamma\right)\left\lVert\hat{G}^{d}(\tilde{\mathrm{x}}_{\text{\tiny AV}},0)\right\rVert^{2}+2\gamma L_{\theta}\left\lVert\tilde{\mathrm{x}}_{\text{\tiny AV}}\right\rVert\left\lVert\tilde{\theta}_{\text{\tiny AV}}\right\rVert+2\gamma^{2}L_{\theta}^{2}\left\lVert\tilde{\theta}_{\text{\tiny AV}}\right\rVert^{2}.$ (55)
@@
$\Delta W(\tilde{\theta}_{\text{\tiny AV}},\tilde{\mathrm{x}}_{\text{\tiny AV}})\!\leq\!-\!\begin{bmatrix}\left\lVert\tilde{\theta}_{\text{\tiny AV}}\right\rVert\\ \left\lVert\tilde{\mathrm{x}}_{\text{\tiny AV}}\right\rVert\end{bmatrix}^{\top}\!\!\!\underbrace{\begin{bmatrix}\alpha\gamma\beta_{\theta}-2\gamma^{2}L_{\theta}^{2}&-\gamma L_{\theta}\\ -\gamma L_{\theta}&\gamma\beta_{x}\end{bmatrix}}_{=:P_{W}(\alpha)}\!\!\begin{bmatrix}\left\lVert\tilde{\theta}_{\text{\tiny AV}}\right\rVert\\ \left\lVert\tilde{\mathrm{x}}_{\text{\tiny AV}}\right\rVert\end{bmatrix}\!$. (56)
Let us set $\alpha>\bar{\alpha}\coloneq(L_{\theta}^{2}\left(1+2\gamma\beta_{x}\right))/\beta_{\theta}\beta_{x}$. By Sylvester Criterion, we get $P_{W}(\alpha)>0$ and, thus, we bound (56) as
$\Delta W\left(\chi_{\text{\tiny AV}}\right)\leq-\lambda_{\text{min}}\left(P_{W}(\alpha)\right)\left\lVert\chi_{\text{\tiny AV}}\right\rVert^{2},$
where $\lambda_{\text{min}}(P_{W}(\alpha))>0$ represents the smallest eigenvalue of $P_{W}$ and $\chi_{\text{\tiny AV}}\coloneq\operatorname{col}(\tilde{\theta}_{\text{\tiny AV}},\tilde{\mathrm{x}}_{\text{\tiny AV}})$. Thus, the exponential stability of the origin is proved (see, e.g. [39, Th. 13.11]).

### v3

@@ VI Conclusions
We proposed DELTA, a data-driven distributed algorithm for aggregative optimization in unknown scenarios. DELTA combines an optimization-oriented part with a tracking mechanism and a learning one. The latter takes on local neural networks elaborating cost samples around the current solution estimates to approximate their gradients. We analyzed DELTA by using tools from system theory based on timescale separation and averaging theory. In detail, we theoretically proved linear convergence in a neighborhood of the optimal solution whose size depends on the given accuracy capabilities of the neural networks. Some numerical tests confirmed our findings.
@@ A-A Preliminaries on Averaging Theory
We report [1, Theorem 2.2.2], which is a useful result in the context of averaging theory. Consider the time-varying system
@@
$f_{\text{\tiny AV}}(\chi)\coloneq\lim_{T\to\infty}\frac{1}{T}\sum_{\tau=\bar{k}+1}^{\bar{k}+T}f(\chi,0,\tau),$ (43)
@@
[1, Theorem 2.2.3] Consider system (42) and let Assumptions 789, and 10 hold. If there exists $\bar{\gamma}_{1}>0$ such that, for all $\gamma\in(0,\bar{\gamma}_{1})$, the origin is an exponentially stable equilibrium point of the averaged system (44), then there exists $\bar{\gamma}\in(0,\bar{\gamma}_{1})$ such that, for all $\gamma\in(0,\bar{\gamma})$, the origin is an exponentially stable equilibrium point of system (42). $\blacksquare$
@@ B-B Proof of Lemma 1
We point out that Assumption (3) implies that $R^{\top}AR$ is Schur. Hence, for all $q>0$, there exist $\Pi_{1}=\Pi_{1}^{\top}>0$ and $\Pi_{2}=\Pi_{2}^{\top}>0$ solving the discrete-time Lyapunov equations
@@
$\left(R^{\top}AR\right)^{\top}\Pi_{2}\left(R^{\top}AR\right)-\Pi_{2}$ $=-I_{(N-1)d},$ (46b)
@@
Being $U$ quadratic, the conditions (24a) and (24c) are satisfied. To conclude the proof, we show that (24b) is verified studying the increment of $U$ along the trajectories of system (23), i.e.,
@@
$\|\Delta\hat{G}_{2}(\tilde{\mathrm{x}}\!+\!x^{\star}\!,\tilde{\mathrm{w}}_{\perp},\tilde{\theta}\!+\!\theta^{\star}\!)\|\leq L_{\sigma}\|R\tilde{\mathrm{w}}_{\perp}\|$, (48)
@@
[…]
$\hskip 12.0pt+\left\lVert(R^{\top}M)\!^{\top}\!\Pi_{2}(R^{\top}M)\right\rVert\left\lVert\Delta\hat{G}_{2}(\mathrm{x},\zeta_{1},\tilde{\theta})\right\rVert^{2}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\!-q\left\lVert\zeta_{1}\right\rVert^{2}-\left\lVert\zeta_{2}\right\rVert^{2}+2L_{\sigma}\left\lVert R^{\top}AR\Pi_{2}RA\right\rVert\left\lVert R\zeta_{1}\right\rVert\left\lVert\zeta_{2}\right\rVert$
$\hskip 12.0pt+\left\lVert(R^{\top}M)\!^{\top}\!\Pi_{2}(R^{\top}M)\right\rVert L_{\sigma}^{2}\left\lVert R\zeta_{1}\right\rVert^{2},$ (49)
where in $(a)$ we use the Cauchy-Schwarz inequality and the definition of $g_{\textsc{b}}$ (cf. (22)), in $(b)$ we use equation (48). Let us define $\kappa_{1}\coloneqq\left\lVert(R^{\top}M)\!^{\top}\!\Pi_{2}(R^{\top}M)\right\rVert L_{\sigma}^{2}\left\lVert R\right\rVert^{2}$ and $\kappa_{12}\coloneq L_{\sigma}\left\lVert R^{\top}AR\Pi_{2}RA\right\rVert\left\lVert R\right\rVert$. Then, we compactly rewrite (49) as
@@
$q>\kappa_{1}\qquad\text{and}\qquad q>\kappa_{1}+\kappa_{12}^{2}.$ (51)
We set $q>\bar{q}_{1}\coloneq\kappa_{1}+\kappa_{12}^{2}$ and, thus, we bound (50) as
@@ C-C Proof of Lemma 2
$\Delta W_{\theta}(\tilde{\theta}_{\text{\tiny AV}},\tilde{\mathrm{x}}_{\text{\tiny AV}})$ $\coloneq\|\tilde{\theta}_{\text{\tiny AV}}-\gamma\nabla_{2}\tilde{\ell}_{\mathcal{K}}(\tilde{\mathrm{x}}_{\text{\tiny AV}},\tilde{\theta}_{\text{\tiny AV}})\|^{2}-\|\tilde{\theta}_{\text{\tiny AV}}\|^{2}$
$\Delta W_{\mathrm{x}}(\tilde{\theta}_{\text{\tiny AV}},\tilde{\mathrm{x}}_{\text{\tiny AV}})$ $\coloneq\|\tilde{\mathrm{x}}_{\text{\tiny AV}}-\gamma\hat{G}^{d}(\tilde{\mathrm{x}}_{\text{\tiny AV}},\tilde{\theta}_{\text{\tiny AV}})\|^{2}-\|\tilde{\mathrm{x}}_{\text{\tiny AV}}\|^{2}.$
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}-\gamma\tfrac{2\mu_{\ell}L_{\ell}}{\mu_{\ell}\!+\!L_{\ell}}\left\lVert\tilde{\theta}_{\text{\tiny AV}}\right\rVert^{2}\!-\!\gamma(\tfrac{2}{\mu_{\ell}\!+\!L_{\ell}}\!-\!\gamma)\left\lVert\nabla_{2}\tilde{\ell}_{\mathcal{K}}(\tilde{\mathrm{x}}_{\text{\tiny AV}},\tilde{\theta}_{\text{\tiny AV}})\right\rVert^{2},$ (55)
where in $(a)$ we use $\nabla_{2}\tilde{\ell}_{\mathcal{K}}(\tilde{\mathrm{x}}_{\text{\tiny AV}},0)=0$, the $\mu_{\ell}$-strong convexity of $\tilde{\ell}_{\mathcal{K}}$ in $\Theta$, and the $L_{\ell}$-Lipschitz continuity of $\nabla_{2}\tilde{\ell}_{\mathcal{K}}$ (cf. Assumption 6). Then, we expand $\Delta W_{\mathrm{x}}$ and obtain
@@
$\!\stackrel{{\scriptstyle(a)}}{{=}}-2\gamma\nabla f_{\sigma}(\tilde{\mathrm{x}}_{\text{\tiny AV}}+x^{\star})^{\top}\tilde{\mathrm{x}}_{\text{\tiny AV}}$
$\hskip 12.0pt-2\gamma(\hat{G}^{d}(\tilde{\mathrm{x}}_{\text{\tiny AV}},\tilde{\theta}_{\text{\tiny AV}})\!-\!\hat{G}^{d}(\tilde{\mathrm{x}}_{\text{\tiny AV}},0))^{\top}\!\tilde{\mathrm{x}}_{\text{\tiny AV}}$
@@
$\!\stackrel{{\scriptstyle(b)}}{{=}}-2\gamma\nabla f_{\sigma}(\tilde{\mathrm{x}}_{\text{\tiny AV}}+x^{\star})^{\top}\tilde{\mathrm{x}}_{\text{\tiny AV}}+2\gamma^{2}\left\lVert\hat{G}^{d}(\tilde{\mathrm{x}}_{\text{\tiny AV}},0)\right\rVert^{2}\!$
$\hskip 13.0pt\!+2\gamma\!\left\lVert\hat{G}^{d}(\tilde{\mathrm{x}}_{\text{\tiny AV}},\tilde{\theta}_{\text{\tiny AV}})\!-\!\hat{G}^{d}(\tilde{\mathrm{x}}_{\text{\tiny AV}},0)\right\rVert\left\lVert\tilde{\mathrm{x}}_{\text{\tiny AV}}\right\rVert$
@@
where in $(a)$ we add $\pm\nabla f_{\sigma}(\tilde{\mathrm{x}}_{\text{\tiny AV}}+x^{\star})\!=\!\hat{G}^{d}(\tilde{\mathrm{x}}_{\text{\tiny AV}},0)$, and $(b)$ uses the triangle and Cauchy-Schwarz inequalities and $\nabla f_{\sigma}(x^{\star})\!=\!\hat{G}^{d}(0,0)\!=\!0$. By Lipschitz continuity of $\nabla_{1}\hat{f}_{i}$, $\nabla_{2}\hat{f}_{i}$, and $\phi_{i}$ (cf. Assumptions 2 and 4), there exists $L_{\theta}\!>\!0$ such that
$\|\hat{G}^{d}(\tilde{\mathrm{x}}_{\text{\tiny AV}},\tilde{\theta}_{\text{\tiny AV}})-\hat{G}^{d}(\tilde{\mathrm{x}}_{\text{\tiny AV}},\bar{\theta}_{\text{\tiny AV}})\|\leq L_{\theta}\|\tilde{\theta}_{\text{\tiny AV}}-\bar{\theta}_{\text{\tiny AV}}\|,$ (57)
for all $\tilde{\mathrm{x}}_{\text{\tiny AV}}\!\in\!\mathbb{R}^{n}$ and $\tilde{\theta}_{\text{\tiny AV}},\bar{\theta}_{\text{\tiny AV}}\!\in\!\mathbb{R}^{m}$. By (57), $\mu_{f}$-strong convexity of $f_{\sigma}$, and $L_{x}$-Lipschitz continuity of $\nabla f_{\sigma}$, we bound (56) as
$\Delta W_{\mathrm{x}}(\tilde{\theta}_{\text{\tiny AV}},\tilde{\mathrm{x}}_{\text{\tiny AV}})$ $\leq-\gamma\tfrac{2\mu_{f}L_{x}}{\mu_{f}+L_{x}}\left\lVert\tilde{\mathrm{x}}_{\text{\tiny AV}}\right\rVert^{2}-\gamma(\tfrac{2}{\mu_{f}+L_{x}}-\gamma)\|\hat{G}^{d}(\tilde{\mathrm{x}}_{\text{\tiny AV}},0)\|^{2}$
$\hskip 11.0pt+2\gamma L_{\theta}\left\lVert\tilde{\mathrm{x}}_{\text{\tiny AV}}\right\rVert\|\tilde{\theta}_{\text{\tiny AV}}\|+2\gamma^{2}L_{\theta}^{2}\|\tilde{\theta}_{\text{\tiny AV}}\|^{2}.$ (58)
@@
$\Delta W(\tilde{\theta}_{\text{\tiny AV}},\tilde{\mathrm{x}}_{\text{\tiny AV}})\!\leq\!-\!\begin{bmatrix}\left\lVert\tilde{\theta}_{\text{\tiny AV}}\right\rVert\\ \left\lVert\tilde{\mathrm{x}}_{\text{\tiny AV}}\right\rVert\end{bmatrix}^{\top}\!\!\!\underbrace{\begin{bmatrix}\alpha\gamma\beta_{\theta}-2\gamma^{2}L_{\theta}^{2}&-\gamma L_{\theta}\\ -\gamma L_{\theta}&\gamma\beta_{x}\end{bmatrix}}_{=:P_{W}(\alpha)}\!\!\begin{bmatrix}\left\lVert\tilde{\theta}_{\text{\tiny AV}}\right\rVert\\ \left\lVert\tilde{\mathrm{x}}_{\text{\tiny AV}}\right\rVert\end{bmatrix}\!$.
Choose $\alpha\!>\!\bar{\alpha}\!\coloneq\!(L_{\theta}^{2}\left(1\!+\!2\gamma\beta_{x}\right))/\beta_{\theta}\beta_{x}$. By Sylvester Criterion, we get $P_{W}(\alpha)>0$ and, thus, we bound $\Delta W(\chi_{\text{\tiny AV}})$ as
$\Delta W\left(\chi_{\text{\tiny AV}}\right)\leq-\lambda_{\text{min}}\left(P_{W}(\alpha)\right)\left\lVert\chi_{\text{\tiny AV}}\right\rVert^{2}$
where $\lambda_{\text{min}}(P_{W}(\alpha))>0$ represents the smallest eigenvalue of $P_{W}$ and $\chi_{\text{\tiny AV}}\coloneq\operatorname{col}(\tilde{\theta}_{\text{\tiny AV}},\tilde{\mathrm{x}}_{\text{\tiny AV}})$. Thus, the exponential stability of the origin is proved (see, e.g. [19, Th. 13.11]).